# SReferences

### v1

@@ Abstract
In today’s digital landscape, journalists urgently require tools to verify the authenticity of facial images and videos depicting specific public figures before incorporating them into news stories. Existing deepfake detectors are not optimized for this detection task when an image is associated with a specific and identifiable individual. This study focuses on the deepfake detection of facial images of individual public figures. We propose to condition the proposed detector on the identity of the identified individual given the advantages revealed by our theory-driven simulations. While most detectors in the literature rely on perceptible or imperceptible artifacts present in deepfake facial images, we demonstrate that the detection performance can be improved by exploiting the idempotency property of neural networks. In our approach, the training process involves double neural-network operations where we pass an authentic image through a deepfake simulating network twice. Experimental results show that the proposed method improves the area under the curve (AUC) from 0.92 to 0.94 and reduces its standard deviation by 17%. For evaluating the detection performance of individual public figures, a facial image dataset with individuals’ names is required, a criterion not met by the current deepfake datasets. To address this, we curated a dataset comprising 32k images featuring 45 public figures, which we intend to release to the public after the paper is published.
1 Introduction
Figure 1: The inference pipeline of the proposed individualized deepfake detector leveraging the near-idempotence property and identity conditioning. The identity conditioning is achieved by combining the identity-aware processing trace and the input identity vector. To leverage the idempotence property, the test image is passed through a reconstruction operator $R$. If the test image exhibits a marginal change in the observed amount of processing traces, the test image is considered “deepfake”; if a significant change is observed, the image is considered “authentic”.
@@ 1 Introduction
A deepfake refers to a seemingly authentic image or video generated by a deep neural network. When it comes to human faces, a manipulation method may comprise reenactment, replacement, editing, and synthesis [41]. While deepfakes can facilitate numerous appealing and advantageous applications, the act of replacing the face in a staged image or video with the face of a public figure can pose a serious threat to the society. Given the continuous influx of deepfake videos on public platforms, journalists need to pay special attention to those that relate to significant public interest, such as those featuring celebrities or politicians [41, 26]. The deepfake generation methods evolved with autoencoder-based approaches [1], GANs [6], and diffusion models [47]. The latest diffusion-based models such as [47, 49] can surpass GAN-based models in producing photorealistic images. Nevertheless, even in the present day, autoencoder-based models remain threatening in terms of malicious use. This is due to the availability of several free, downloadable, and user-friendly applications built on autoencoder, such as FaceSwap [5], Faceswap-GAN [6], DeepFaceLab [4], and df [2]. In this work, we focus on Faceswap-GAN.
@@
Our approach to deepfake detection draws inspiration from a series of studies leveraging the near-idempotence property of an operation. This method has been particularly effective in various image forensics tasks, including double JPEG compression detection, unknown video codec identification, and source camera identification  [32, 53, 9, 23, 40]. In these studies, researchers leverage the near-idempotence of a respective operation, such as certain type of JPEG compression, video compression, or color demosaicing algorithm. The strict idempotence property asserts that an idempotent operation, $f(\cdot)$, results in no change to $f(x)$ when it is applied iteratively, i.e., $f(f(x))=f(x)$. Using slightly different terminology, if $f(f(x))$ approximately equals $f(x)$, the operation is nearly idempotent. In many detection problems of multimedia forensics, the nearly idempotent nature of a forgery method allows an analyst to apply the forgery operation multiple times and observe the changes to determine whether the input was forged for the first time, i.e., input forged for more than once will exhibit minimal changes.
In this work, we demonstrate that near-idempotence is also applicable to the neural network-based Faceswap-GAN [6]. To explore this, we emulate the potential deepfake operation that an attacker might employ, utilizing publicly available data of a public figure and making assumptions about the neural network architecture. Fig. 1 illustrates the inference pipeline of the proposed detector. We feed a test image into the emulated deepfake generator. The expected change in the image due to this operation is dependent on whether the image has undergone a similar operation before. If the image is a deepfake, the near-idempotence property ensures that the change will be minimal. From the standpoint of the deepfake feature extractor, a deepfake image will exhibit processing traces both before and after the operation, leading to subtle observed changes. Conversely, an authentic image without the deepfake operation lacks any processing traces of the neural network, resulting in a significant observable change. The contributions of this paper are threefold.
@@
2 Related Work
2.1 Generation of Deepfake Faces
@@
Early methods of face-swapping such as Bitouk et al. [10] were limited to using two images of two particular persons with similar poses. The images were first aligned with the help of landmark detection, then cropped, and postprocessed including color correction. Subsequent researchers [19] improved those with a 3-D facial model from the source video. The next advancement emerged after the proposal of a deep-learning-based face-swapping architecture [1] built upon one shared encoder and two individual decoders. Faceswap-GAN [6] is the GAN improvement over [1] where the performance of shared encoder and individual decoders further improve as a result of the GAN’s internal interplay mechanism between the generator and discriminator. However, the architectures of [1, 6] can swap faces between only two identities involved in training. Researchers have proposed identity agnostic architectures decoupling the identity extraction from the attribute extraction [8, 45, 42, 43, 37].
2.2 Protection Against Deepfake
@@
Researchers have been exploring different methods to detect deepfakes. In the first category, the artifacts of synthetic videos are exploited for deepfake detection such as the absence of eye blinking [38], inconsistency in head pose. [54], disparities in color components [36], and inconsistency between inner face and outer face [25]. In the second category, researchers used either an end-to-end convolutional neural network (CNN) structure  [51] or a combined CNN with a recurrent neural network (RNN) [29]. In the third category, researchers exploit processing traces left by the neural networks for deepfake detection. The researchers exploited the features like spatial domain local convolutional features [28], spectral distortion caused by up-convolutions [26], and upsampling artifacts in the frequency domain [27].
@@
Identity-driven Deepfake Detection. Instead of detecting deepfake videos for the whole population, recent work also exploited characteristics of a specific person for deepfake detection. Agarwal et al. [7] targeted deepfake videos of a specific individual by capturing speaking patterns. Cozzolino et al. [21] proposed to learn the temporal features of how a specific person moves and talks. Dong et al. [25] calculated $\ell^{2}$ distance between the computed identity vector from the inner face and the expected identity vector drawn from a reference set of identity vectors. In this work, we extract the deepfake traces conditioned on the identity.
2.3 Idempotency as a Multimedia Forensics Tool
@@
In multimedia forensics, one way to detect counterfeiting is to exploit the near-idempotence property, i.e., the minor changes caused by the repetitive application of adversarial operations. It shares the same sprit of the law of diminishing returns, a widely used concept in economics [13, 50]. The detection of double JPEG compression, source camera identification, and video codec identification are three exemplary applications of the near-idempotence property. The ratio of stable image blocks has been used by researchers to detect the number of prior JPEG compressions [35, 15]. Huang et al. [32] found that the number of dissimilar JPEG coefficients between two subsequent JPEG compression decreases monotonically. Bestagini et al. [9] detected unknown video encoding by recompressing a video with each of the candidates. For source camera identification, the researchers have leveraged the near-idempotence property of an auto-white balancing method [23] and that of color demosaicing strategy [40]. In economics, the law of diminishing returns states that additional inputs to a fixed amount of identical inputs increase productivity at a decreasing rate [13]. If the additional inputs are considered repetitive operations, then the law of diminishing returns may be considered as near-idempotence. In this study, we show that the near-idempotence property of neural networks assists in deepfake image detection.
2.4 Unsupervised Pretraining
@@
Unsupervised pertaining has been proposed for feature extraction for many tasks of computer vision. Chen et al. [18] found that larger networks, for example, larger ResNet, pretrained in an unsupervised manner followed by supervised training with only $10\%$ of labeled data can outperform fully supervised networks for general computer vision tasks. Newell and Deng [44] showed that pretrained networks are more advantageous in low data regimes compared to ubiquitous data. Their results suggest that pretrained networks should be tested on diverse downstream tasks. Bulat et al. [14] proposed task-agnostic self-supervised pretraining on in-the-wild facial data for representation learning. Zheng et al. [58] proposed weakly supervised facial representation learning using vast facial images available on the web with linguistic descriptions. In this work, we utilize the facial features from Bulat et al. [14] to additionally learn the deepfake traces.
3 Threat Model
@@
In this work, we consider an attacker who is smart enough to find and use open-source face-swapping software such as [6, 5, 1] on the facial images from the publicly available videos of a public figure. More specifically, we consider Faceswap-GAN [6] as a potential method that the attacker can use. The attacker is free to use any public or private videos of a second person to depict a story and they want to convince the public of the involvement of a targeted public figure. For example, the attacker can record prearranged videos at a professional studio and later replace the actor’s face with that of the public figure. The attacker can harvest videos of the public figure from multiple sources, including social media, news channels, movies, and YouTube. Different sources of videos offer varied image quality, compression levels, and processing histories. For example, public interview videos of a public figure available on YouTube are expected to be less edited than video clips from movies. In our proposed detection method, we assume that we, as forensic analysts, have access to the various sources of public figure videos, but we do not know exactly from what source the attacker took videos for deepfake generation. For example, the attacker can use videos from social media, where we will only use public interview recordings of that public figure to train the neural network based detector.
4 Proposed Detector via Near-Idempotence and Identity Conditioning
@@
In the challenge of identifying deepfake faces for public figures, we confront an image of unknown authenticity, claimed to be a specific public figure. Our approach to addressing this problem makes use of the extensive collection of authentic images or videos of the said public figure from YouTube. The training process of our proposed deepfake detector is depicted in Fig. 2 and the inference pipeline is shown in Fig. 1.
@@
Figure 2: The training pipeline of the proposed deepfake detector leveraging the near-idempotence property of the deepfake generator. A side-by-side comparison with conventional deepfake detectors is also shown. In the proposed method, an authentic image is passed through a deepfake simulating network or reconstruction operator, twice. Due to the near-idempotence property, the features for the first and the second outputs will be nearly identical. The features are obtained from an identity-aware feature extractor that is trained separately. We freeze the feature extractor network and train a Siamese network and an identity decoder to increase the Euclidean distance between the first pair (consisting of the authentic image and the first output image) and to decrease the Euclidean distance between the second pair (consisting of the first and the second output images).
@@
Our proposed detector has four distinct components. First, the reconstruction operator is a neural network operation that stimulates the deepfake generation operation for a public figure. We found this operation nearly idempotent. Second, the feature extractor is finetuned with a teacher network and is able to capture the identity information while extracting the features. Third, the identity decoder takes as input the explicit identity, i.e., the index of the public figure, and learns as a constant identity vector that arguments the feature space. It contains the necessary person-specific information of that public figure, and when combined with the identity-aware feature can effectively compute the deepfake features conditioned on identity. Fourth, the Siamese network serves as the ultimate binary classification block in the proposed architecture. It learns to extract the features linked to the idempotency of the deepfake operation. It produces a larger distance before and after reconstruction for a test authentic image and a smaller distance for a test deepfake image.
4.1 Reconstruction Operator and Idempotence-Driven Detection
@@
We employ a dedicated reconstruction operator $R$ for each public figure as shown in Fig. 1 and  Fig. 2. When the original image is authentic, the first operation generates a deepfake image, and the second operation produces a doubly processed deepfake. We verified experimentally that the reconstruction operator $R$ serves as a reliable approximation of a specific type of deepfake generation tool, such as FaceSwap-GAN [6], and that the deepfake generation process is nearly idempotent. In this context, the distance between a deepfake image and its corresponding doubly processed deepfake tends to be close to zero. This characteristic is leveraged in the training and inference system.
@@
The reconstruction operator $R$ exhibits near-zero changes to a deepfake image due to the near-idempotence. Consequently, the feature level Euclidean distance between the two is expected to be small. On the other hand, an authentic image and its corresponding processed image will be substantially different as the operation leaves discernible traces in the processed image. Considering the capability of our deepfake feature extractor (see Sec. 4.2) to detect these traces, the features will exhibit significant dissimilarity, resulting in a higher distance compared to the deepfake scenario.
Based on the above considerations, the initial problem of detecting whether an image is authentic or deepfake is now reframed as evaluating the change of the image in the feature space through the reconstruction operation. When this change, quantified as the Euclidean distance, approaches zero, the image is classified as a deepfake; otherwise, it is considered authentic. Denoting the input image by f, the reframed problem is to evaluate whether f and $R(\text{f})$ are the same or not, where $R$ is our reconstruction operator. Treating f and $R(\text{f})$ as two inputs, we note that the Siamese network [12] is a powerful approach for discerning similarity or dissimilarity between two inputs. Our use of the Siamese network will be discussed in Sec. 4.4.
4.2 Identity-Aware Feature Extractor
Motivation. Conventional deepfake feature extraction network $B(\cdot)$ extracts the deepfake features $B({\rm f})$ for a test image ${\rm f}$ ignoring the person identity ${\rm I}$ [11, 22, 57] or considers the identity features irrelevant to forgery detection [30, 52]. Our work found that the identity-aware feature, $B^{{}^{\prime}}({\rm f})$, which extracts identity information in addition to the deepfake features, is more effective for deepfake detection. This may be explained that a distinct extracted feature may not be equally distinguishable for every identity for the classification. If a feature extractor does not allow the passing of the identity information, the later network can not learn the statistics of the features individually for each identity. This will be limited to learning the average pattern. Such average distributions of the features will lead to the error probability of the Bayesian classifier as follows:
$P_{\text{e}}^{\text{com}}\!=\!\mathbb{P}(H_{0})\mathbb{P}(C\!=\!1|H_{0})\!+\!\mathbb{P}(H_{1})\mathbb{P}(C\!=\!0|H_{1}).$ (1)

### v2

# Individualized Deepfake Detection Exploiting Traces Due to Double Neural-Network Operations

@@ Abstract
In today’s digital landscape, journalists urgently require tools to verify the authenticity of facial images and videos depicting specific public figures before incorporating them into news stories. Existing deepfake detectors are not optimized for this detection task when an image is associated with a specific and identifiable individual. This study focuses on the deepfake detection of facial images of individual public figures. We propose to condition the proposed detector on the identity of an identified individual, given the advantages revealed by our theory-driven simulations. While most detectors in the literature rely on perceptible or imperceptible artifacts present in deepfake facial images, we demonstrate that the detection performance can be improved by exploiting the idempotency property of neural networks. In our approach, the training process involves double neural-network operations where we pass an authentic image through a deepfake simulating network twice. Experimental results show that the proposed method improves the area under the curve (AUC) from 0.92 to 0.94 and reduces its standard deviation by 17%. To address the need for evaluating detection performance for individual public figures, we curated and publicly released a dataset of $\sim$32k images featuring 45 public figures, as existing deepfake datasets do not meet this criterion.
Index Terms:
Deepfake detection, double operations, double JPEG compression, Siamese neural network, manifold learning.
I Introduction
Figure 1: (a) Comparison of AUC performance for nine off-the-shelf deepfake detection methods (listed in TABLE I), two fine-tuned methods, and the proposed method, evaluated on GAN-based Faceswap-GAN [1] deepfakes and diffusion-based DiffSwap [2] deepfakes. Square markers denote methods without finetuning (unfilled) and with finetuning (filled), while the star marker highlights the proposed method. Method names are labeled near their respective markers for better visualization. (b) The inference pipeline of the proposed individualized deepfake detector leveraging the near-idempotence property and identity conditioning. The identity conditioning is achieved by combining the identity-aware processing trace and the input identity vector. To leverage the idempotence property, the test image is passed through a reconstruction operator $R$. If the test image exhibits a marginal change in the observed amount of processing traces, the test image is considered “deepfake”; if a significant change is observed, the image is considered “authentic.” (c) t-SNE visualization of authentic, deepfake, and doubly-processed images and their corresponding vector shifts in t-SNE feature space across the two transformations. Red arrows indicate the vector shifts for the first transformation, while black arrows represent shifts during the second transformation. The first transformation causes a significant change when a deepfake operator initially processes an authentic image. In contrast, the second transformation results in only minor shifts. For more details, please refer to Section V-F.
@@ I Introduction
A deepfake refers to a seemingly authentic image or video generated by a deep neural network. When it comes to human faces, a manipulation method may comprise reenactment, replacement, editing, and synthesis [3]. While deepfakes can facilitate numerous appealing and advantageous applications, the act of replacing the face in a staged image or video with the face of a public figure can pose a serious threat to society. Given the continuous influx of deepfake videos on public platforms, journalists need to pay special attention to those that relate to significant public interest, such as those featuring celebrities or politicians [3, 4]. The deepfake generation methods evolved with autoencoder-based approaches [5], GANs [1], and diffusion models [6]. The latest diffusion-based models, such as [2, 6, 7], can surpass GAN-based models in producing photorealistic images. Nevertheless, even in the present day, autoencoder-based models remain threatening in terms of malicious use. This is due to the availability of several free, downloadable, and user-friendly applications built on autoencoder, such as FaceSwap [8], Faceswap-GAN [1], DeepFaceLab [9], and df [10]. In this work, we focus on GAN-based Faceswap-GAN [1] and diffusion-based DiffSwap [2].
@@
Our approach to deepfake detection draws inspiration from a series of studies leveraging the near-idempotence property of an operation. This method has been particularly effective in various image forensics tasks, including double JPEG compression detection, unknown video codec identification, and source camera identification  [11, 12, 13, 14, 15]. In these studies, researchers leverage the near-idempotence of a respective operation, such as certain type of JPEG compression, video compression, or color demosaicing algorithm. The strict idempotence property asserts that an idempotent operation, $g(\cdot)$, results in no change to $g(x)$ when it is applied iteratively, i.e., $g(g(x))=g(x)$. Using slightly different terminology, if $g(g(x))$ approximately equals $g(x)$, the operation is nearly idempotent. In many detection problems of multimedia forensics, the nearly idempotent nature of a forgery method allows an analyst to apply the forgery operation multiple times and observe the changes to determine whether the input was forged for the first time, i.e., input forged for more than once will exhibit minimal changes.
In this work, we demonstrate that near-idempotence is also applicable to the neural network-based Faceswap-GAN [1] and DiffSwap [2], as demonstrated in Fig. 1(c). To explore this, we emulate a potential deepfake operation that an attacker might employ, utilizing publicly available data of a public figure and employing a neural network architecture to replicate the functionality of a deepfake generation tool. Referring to it as the reconstruction operator $R$, Fig. 1(b) illustrates the inference pipeline of the proposed detector. We feed a test image into the emulated deepfake generator. The expected change in the image due to this operation is dependent on whether the image has undergone a similar operation before. If the image is a deepfake, the near-idempotence property ensures that the change will be minimal. From the standpoint of the deepfake feature extractor, a deepfake image will exhibit processing traces both before and after the operation, leading to subtle observed changes. Conversely, an authentic image without the deepfake operation lacks any processing traces of the neural network, resulting in a significant observable change. The contributions of this paper are threefold.
@@
The remainder of this paper is organized as follows. Section II discusses the existing literature on deepfake generation, detection, and approaches related to the proposed method. Section III introduces the threat model, while Section IV presents the proposed deepfake detection method based on near-idempotence and identity conditioning. Section V showcases the experimental results, followed by Section VI, highlighting the key findings of this work. Finally, Section VII concludes the paper.
II Related Work
II-A Generation of Deepfake Faces
@@
Early methods of face-swapping, such as Bitouk et al. [16], were limited to using two images of two particular persons with similar poses. The images were first aligned with the help of landmark detection, then cropped and postprocessed, including color correction. Subsequent researchers [17] improved those with a 3-D facial model from the source video. The next advancement emerged after the proposal of a deep-learning-based face-swapping architecture [5] built upon one shared encoder and two individual decoders. Faceswap-GAN [1] is the GAN improvement over faceswap [5], where the performance of shared encoder and individual decoders further improve as a result of the GAN’s internal interplay mechanism between the generator and discriminator. However, the architectures proposed in [5, 1] can only swap faces between the two identities involved in training. Researchers have proposed identity-agnostic architectures that decouple identity extraction from attribute extraction [18, 19, 20, 21, 22]. Recent works [23, 24] demonstrate that denoising diffusion models (DDMs) can significantly reduce the performance of deepfake detectors trained on images not generated by DDMs. For example, DiffSwap [2] considers face swapping as a conditional image inpainting task, where the denoising network is conditioned on the identity features of the source image and the facial landmarks of the target image. DiffFace [25] employs a diffusion model with a facial guidance mechanism incorporating three distinct control components for identity, semantic features, and gaze for maintaining consistent pose and facial attributes.
II-B Protection Against Deepfakes
@@
Researchers have explored a variety of techniques for deepfake detection. Some exploit the artifacts of synthetic videos, such as the absence of eye blinking [26], inconsistency in head pose [27], disparities in color components [28], and inconsistency between inner face and outer face [29]. Some other researchers opt for a complete data-driven approach by using either an end-to-end convolutional neural network (CNN) structure [30] or a combined CNN with a recurrent neural network (RNN) [31]. Researchers have also exploited processing traces left by the neural networks for deepfake detection. The researchers exploited the features such as spatial domain local convolutional features [32, 33, 34, 35, 36, 37, 38] and spectral distortion or upsampling artifacts in the frequency domain [4, 39, 40, 41].
@@
Instead of detecting deepfake videos for the whole population, the characteristics of a specific person have also been exploited. Agarwal et al. [42] targeted deepfake videos of a specific individual by capturing speaking patterns. Cozzolino et al. [43] proposed to learn the temporal features of how a specific person moves and talks. Dong et al. [29] calculated $\ell^{2}$ distance between the computed identity vector from the inner face and the expected identity vector drawn from a reference set of identity vectors. In this work, we extract the deepfake traces conditioned on the identity.
II-C Idempotency as a Multimedia Forensics Tool
@@
In multimedia forensics, one way to detect counterfeiting is to exploit the near-idempotence property, i.e., the minor changes caused by the repetitive application of adversarial operations. It shares the same spirit of the law of diminishing returns, a widely used concept in economics [44, 45]. The detection of double JPEG compression, source camera identification, and video codec identification are three exemplary applications of the near-idempotence property. The ratio of stable image blocks has been used by researchers to detect the number of prior JPEG compressions [46, 47]. Huang et al. [11] found that the number of dissimilar JPEG coefficients between two subsequent JPEG compression decreases monotonically. Bestagini et al. [13] detected unknown video encoding by recompressing a video with each of the candidates. For source camera identification, the researchers have leveraged the near-idempotence property of an auto-white balancing method [14] and that of color demosaicing strategy [15]. In economics, the law of diminishing returns states that additional inputs to a fixed amount of identical inputs increase productivity at a decreasing rate [44]. If the additional inputs are considered repetitive operations, then the law of diminishing returns may be regarded as near-idempotence. In this study, we show that the near-idempotence property of neural networks assists in deepfake image detection.
II-D Unsupervised Pretraining
@@
Unsupervised pretaining has been proposed for feature extraction for many tasks of computer vision. Chen et al. [48] found that larger networks, for example, larger ResNet, pretrained in an unsupervised manner followed by supervised training with only $10\%$ of labeled data can outperform fully supervised networks for general computer vision tasks. Newell and Deng [49] showed that pretrained networks are more advantageous in low data regimes compared to ubiquitous data. Their results suggest that pretrained networks should be tested on diverse downstream tasks. Bulat et al. [50] proposed task-agnostic self-supervised pretraining on in-the-wild facial data for representation learning. Zheng et al. [51] proposed weakly supervised facial representation learning using vast facial images available on the web with linguistic descriptions. In this work, we fine-tune the facial features from Bulat et al. [50] to learn the deepfake traces.
III Threat Model
@@
In this work, we consider an attacker who is smart enough to find and use open-source face-swapping software such as [1, 8, 5, 2] on the facial images from the publicly available videos of a public figure. More specifically, we consider Faceswap-GAN [1] and DiffSwap [2] as potential methods that the attacker can use. The attacker is free to use any public or private videos of a second person to depict a story and convince the public of the involvement of a targeted public figure. For example, the attacker can record prearranged videos at a professional studio and later replace the actor’s face with that of a public figure. The attacker can harvest videos of the public figure from multiple sources, including social media, news channels, movies, and YouTube. Different sources of videos offer varied image quality, compression levels, and processing histories. For example, public interview videos of a public figure available on YouTube are expected to be less edited than video clips from movies. In our proposed detection method, we assume that we, as forensic analysts, have access to the various sources of public figure videos, but we do not know exactly from what source the attacker took videos for deepfake generation. For example, the attacker can use videos from social media, where we will only use public interview recordings of that public figure to train the neural network-based detector.
IV Proposed Detector via Near-Idempotence and Identity Conditioning
@@
In the challenge of identifying deepfake faces for public figures, we confront an image of unknown authenticity, claimed to be a specific public figure. Our approach to addressing this problem makes use of the extensive collection of authentic images or videos of the said public figure from YouTube. The training process of our proposed deepfake detector is depicted in Fig. 2, and the inference pipeline is shown in Fig. 1(b).
@@
Figure 2: The training pipeline of the proposed deepfake detector leveraging the near-idempotence property of the deepfake generator. A side-by-side comparison with conventional deepfake detectors is also shown. In the proposed method, an authentic image is passed through a deepfake simulating network or reconstruction operator twice. Due to the near-idempotence property, the features for the first and the second outputs will be nearly identical. The features are obtained from an identity-aware feature extractor that is trained separately. We freeze the feature extractor network and train a Siamese network and an identity decoder to increase the Euclidean distance between the first pair (consisting of the authentic image and the first output image) and to decrease the Euclidean distance between the second pair (consisting of the first and the second output images).
@@
Our proposed detector has four distinct components. First, the reconstruction operator is a neural network operation that stimulates the deepfake generation operation for a public figure. We found this operation nearly idempotent. Second, the feature extractor is finetuned with a teacher network and is able to capture the identity information while extracting the features. Third, the identity decoder takes as input the explicit identity, i.e., the index of the public figure, and learns as a constant identity vector that arguments the feature space. It contains the necessary person-specific information of that public figure and, when combined with the identity-aware feature, can effectively compute the deepfake features conditioned on identity. Fourth, the Siamese network serves as the ultimate binary classification block in the proposed architecture. It learns to extract the features linked to the idempotency of the deepfake operation. It produces a larger distance before and after reconstruction for a test authentic image and a smaller distance for a test deepfake image.
IV-A Reconstruction Operator and Idempotence-Driven Detection
@@
We employ a dedicated reconstruction operator $R$ for each public figure as shown in Fig. 1(b) and Fig. 2. When the original image is authentic, the first operation generates a deepfake image, and the second operation produces a doubly processed deepfake. We verified experimentally that the reconstruction operator $R$ serves as a reliable approximation of a specific type of deepfake generation tool, such as FaceSwap-GAN [1], and that the deepfake generation process is nearly idempotent. In this context, the distance between a deepfake image and its corresponding doubly processed deepfake tends to be close to zero. This characteristic is leveraged in the training and inference system.
@@
The reconstruction operator $R$ exhibits near-zero changes to a deepfake image due to the near-idempotence. Consequently, the feature level Euclidean distance between the two is expected to be small. On the other hand, an authentic image and its corresponding processed image will be substantially different as the operation leaves discernible traces in the processed image. Considering the capability of our deepfake feature extractor (see Section IV-B) to detect these traces, the features will exhibit significant dissimilarity, resulting in a higher distance compared to the deepfake scenario.
Based on the above considerations, the initial problem of detecting whether an image is authentic or deepfake is now reframed as evaluating the change of the image in the feature space through the reconstruction operation. When this change, quantified as the Euclidean distance, approaches zero, the image is classified as a deepfake; otherwise, it is considered authentic. Denoting the input image by f, the reframed problem is to evaluate whether f and $R(\text{f})$ are the same or not, where $R$ is our reconstruction operator. Treating f and $R(\text{f})$ as two inputs, we note that the Siamese network [52] is a powerful approach for discerning similarity or dissimilarity between two inputs. Our use of the Siamese network will be discussed in Section IV-D.
IV-B Identity-Aware Feature Extractor
IV-B1 Motivation
Conventional deepfake feature extraction network $B(\cdot)$ extracts the deepfake features $B({\rm f})$ for a test image ${\rm f}$ ignoring the person identity ${\rm I}$ [37, 53, 54] or considers the identity features irrelevant to forgery detection [55, 56]. Our work found that the identity-aware feature, $B^{{}^{\prime}}({\rm f})$, which extracts identity information in addition to the deepfake features, is more effective for deepfake detection. This may be explained by the fact that a distinct extracted feature may not be equally distinguishable for every identity for the classification. If a feature extractor does not allow the passing of the identity information, the later network can not learn the statistics of the features individually for each identity. This will be limited to learning the average pattern. Such average distributions of the features will lead to the error probability of the Bayesian classifier as follows:
$P_{\text{e}}^{\text{com}}=\mathbb{P}(H_{0})\,\mathbb{P}\left(C\!=\!1\mid H_{0}\right)+\mathbb{P}(H_{1})\,\mathbb{P}\left(C\!=\!0\mid H_{1}\right).$ (1)